# Minimum Strict Consistent Subset in Paths, Spiders, Combs and Trees

### v2

@@ Abstract
In a connected simple graph $G=(V,E)$, each vertex of $V$ is colored by a color from the set of colors $C=\{c_{1},c_{2},\dots,c_{\alpha}\}$. We take a subset $S$ of $V$, such that for every vertex $v$ in $V\setminus S$, at least one vertex of the same color is present in its set of nearest neighbors in $S$. We refer to such a $S$ as a consistent subset (CS). The Minimum Consistent Subset (MCS) problem is the computation of a consistent subset of the minimum size. It is established that MCS is NP-complete for general graphs, including planar graphs. We expand our study to interval graphs and circle graphs in an attempt to gain a complete understanding of the computational complexity of the MCS problem across various graph classes. The strict consistent subset is a variant of consistent subset problems. We take a subset $S^{\prime}$ of $V$, such that for every vertex $v$ in $V\setminus S^{\prime}$, all the vertices in its set of nearest neighbors in $S$ have the same color as $v$. We refer to such a $S^{\prime}$ as a strict consistent subset (SCS). The Minimum Strict Consistent Subset (MSCS) problem is the computation of a consistent subset of the minimum size.
We demonstrate that MSCS is NP-hard in general graphs. We show a 2-approximation in trees. Later, we show polynomial-time algorithms in trees. Later, we demonstrate faster polynomial-time algorithms in paths, spiders, and combs.
Keywords Minimum Strict Consistent Subset  $\cdot$ MSCS for General Graphs  $\cdot$ MSCS  $\cdot$ MCS  $\cdot$ MSCS in Paths  $\cdot$ MSCS in Trees  $\cdot$ MSCS in Spiders  $\cdot$ MSCS in Combs  $\cdot$ Paths  $\cdot$ Trees  $\cdot$ Combs  $\cdot$ Spiders  $\cdot$ Nearest-Neighbor Classification  $\cdot$ Strict Consistent Subset  $\cdot$ Approximation Algorithm  $\cdot$ NP-hardness  $\cdot$ Algorithm  $\cdot$ Polynomial-time Algorithm
@@ 1 Introduction
Many supervised learning approaches use a colored training dataset $T$ in a metric space $(X,d)$ as input, with each element $t\in T$ assigned a color $c_{i}$ from the set of colors $C=\{c_{1},c_{2},\dots,c_{\alpha}\}$. The purpose is to preprocess $T$ in order to meet specific optimality requirements by immediately assigning color to any uncolored element in $X$. The closest neighbor rule, which assigns a color to each uncolored element $x$ based on the colors of its $k$ closest neighbors in the training dataset $T$ (where $k$ is a fixed integer), is a popular optimality criterion. The efficiency of such an algorithm is determined by the size of the training dataset. As a result, it is necessary to reduce the size of the training set while retaining the distance data. Hart [1] introduced this concept in $1968$ with the minimum consistent subset (MCS) problem. Given a colored training dataset $T$, the goal is to find a subset $S\subseteq T$ of the minimum cardinality such that the color of every point $t\in T$ is the same as one of its nearest neighbors in $S$. Over $2800$ citations to [1] on Google Scholar demonstrate that the MCS problem has found several uses since its inception. [2] demonstrated that the MCS problem for points in $\Re^{2}$ under the Euclidean norm is NP-complete if at least three colors color the input points. Furthermore, it is NP-complete even with two colors [3]. Recently, it was established that the MCS problem is W[1]-hard when parameterized by the output size [4].[2] and [5] offer algorithms for the MCS problem in $\Re^{2}$.
1.1 Notaions and Definitions
@@
This paper explores the minimum consistent subset problem when $(\mathcal{X},d)$ is a graph metric. Most graph theory symbols and notations are standard and come from [6]. For every graph $G$, we refer to the set of vertices as $V$ or $V(G)$, and the set of edges as $E$ or $E(G)$. Consider a graph $G$ with a vertex coloring function $C:V(G)\rightarrow\{c_{1},c_{2},\dots,c_{\alpha}\}$. Let $C(U)$ represent the set of colors of the vertices in $U$. Formally, $C(U)=\{C(u):u\in U\}$. For any two vertices $u,v\in V(G)$, we use $\mathrm{d}(u,v)$ to signify the shortest path distance (number of edges) between $u$ and $v$ in $G$. $\mathrm{d}(v,U)=\min_{u\in U}\mathrm{d}(v,u)$ for a vertex $v\in V(G)$ and any subset of vertices $U\subseteq V$. For any graph $G$ and any vertex $v\in V(G)$, let $\mathrm{N}(v)=\{u\mbox{ }|\mbox{ }u\in V(G),(u,v)\in E(G)\}$ indicate the set of neighbors of $v$ and $\mathrm{N}[v]=\{v\}\cup\mathrm{N}(v)$. For $U\subset V$, $\mathrm{N}[U]=\cup_{u\in U}\mathrm{N}[u]$. The $r$-th neighborhood of $v$ is denoted by $\mathrm{N}^{r}[v]=\mathrm{N}[\mathrm{N}^{r-1}[v]]$, which is recursively defined. Note that $\mathrm{N}^{1}[v]=\mathrm{N}[v]$ and $\mathrm{N}[\emptyset]=\emptyset$. However, for an induced graph $G_{1}$ of $G$, the set of neighbor vertices of a vertex $v\in V(G)$ in $G_{1}$ is denoted by $\mathrm{N}_{G_{1}}(v)$ and defined by $\mathrm{N}_{G_{1}}(v)=\{u\mbox{ }|\mbox{ }u\in V(G_{1}),(u,v)\in E(G)\}$. $\mathrm{N}_{G_{1}}[v]=\{v\}\cup\mathrm{N}_{G_{1}}(v)$. Also, for $U\subset V(G_{1})$, $\mathrm{N}_{G_{1}}[U]=\cup_{u\in U}\mathrm{N}_{G_{1}}[u]$ and $\mathrm{N}_{G_{1}}^{r}[v]=\mathrm{N}_{G_{1}}[\mathrm{N}_{G_{1}}^{r-1}[v]]$. The distance between two subgraphs $G_{1}$ and $G_{2}$ in $G$ is represented as $\mathrm{d}(G_{1},G_{2})=\min\{\mathrm{d}(v_{1},v_{2})\mbox{ }|\mbox{ }v_{1}\in V(G_{1}),v_{2}\in V(G_{2})\}$. Therefore, $\mathrm{N}^{r}[v]=\{u\in V\mbox{ }|\mbox{ }\mathrm{d}(u,v)\leq r\}$. For any subset of vertices $U\subseteq V$ in a graph $G$, $G[U]$ represents the subgraph of $G$ induced on $U$. The nearest neighbor of $v$ in the set $U$ is indicated as $\mathrm{NN}(v,U)$, which is formally defined as $\{u\in U\mbox{ }|\mbox{ }\mathrm{d}(v,u)=\mathrm{d}(v,U)\}$. Also, for $X,U\subseteq V$, $\mathrm{NN}(X,U)=\cup_{v\in X}\{u\in U\mbox{ }|\mbox{ }\mathrm{d}(v,u)=\mathrm{d}(v,U)\}$. The shortest path in $G$ is $\mathrm{path}(v,u)$, which connects vertices $v$ and $u$. $\mathrm{path}(v,U)=\{\mathrm{path}(v,u)\mbox{ }|\mbox{ }u\in\mathrm{NN}(v,U)\}$.
If there is a vertex $u\in\mathrm{NN}(v,U)$ such that $C(v)=C(u)$, then $v$ is said to be covered by the set $U$. We sometimes use only $u$ covers $v$ where $u\in U$. $\mathrm{COV}(v,U)$ refers to the set of vertices in $U$ that cover $v$. $\mathrm{COV}(v,U)=\{u\in U\mbox{ }|\mbox{ }u\in\mathrm{NN}(v,U)\mbox{ and }C(u)=C(v)\}$, and for a subset $X\subseteq V$, $\mathrm{COV}(X,U)=\cup_{v\in X}\{u\in U\mbox{ }|\mbox{ }u\in\mathrm{NN}(v,U)\mbox{ and }C(u)=C(v)\}$. Thus, $u\in U$ covers $v$ indicates that $u\in\mathrm{COV}(v,U)$. If $\mathrm{COV}(v,U)=\emptyset$, then $v$ is not covered by the set $U$; otherwise, $v$ is covered by the set $U$. Similarly, if $C(X)=C(\mathrm{COV}(X,U))$, then $X$ is covered by $U$. If $v$ is not covered by the set $U$, $C(v)\neq C(\mathrm{COV}(v,U))$. For an induced graph $G(U)$, $\textup{{MCS}}(U)$ denotes the consistent subset of $G(U)$. However, simply MCS and $\textup{{MCS}}(V(G))$ are the same. Without loss of generality, we shall use $[n]$ to denote the set of integers $\{1,\ldots,n\}$. A subset $X$ is said to be a consistent subset (in short CS) for $(G,C)$ if every $v\in V(G)$, $C(v)\in C(\mathrm{NN}(v,X))$. A subset $S$ is said to be a strict consistent subset (in short SCS) for $(G,C)$ if every $v\in V(G)$, $C(v)=C(\mathrm{NN}(v,S))$. The strict consistent subset problem in graphs is defined as follows:
@@
MCS on Graphs Input: A graph $G$, a coloring function $C:V(G)\rightarrow\{c_{1},c_{2},\dots,c_{\alpha}\}$, and an integer $\ell$. Question: Does there exist a strict consistent subset of size $\leq\ell$ for $(G,C)$?
Figure 1(A) shows an example of MCS, MSCS etc. Minimum consistent subset (MCS) refers to a consistent subset with a minimum size. For an induced subgraph $G^{{}^{\prime}}=G(U)$ with $U\subseteq V$, the minimum consistent subset of $G^{{}^{\prime}}$ is denoted by $\textup{{MCS}}(G^{{}^{\prime}})$ or $\textup{{MCS}}(U)$ and a consistent subset is denoted by $\mathcal{S}(G^{{}^{\prime}})$ or $\mathcal{S}(U)$.
@@
Figure 1: (A) An MCS for a two-colored unweighted tree. (B) Blocks in a tree. and (C) an MCSS for the same tree.
@@ 2.1 Previous Results
Banerjee et al. [7] demonstrated that the MCS is  with when parameterized by the size of the minimum consistent set, even with only two colors. Under basic complexity-theoretic assumptions for generic networks, a FPT algorithm parameterized by $(c+\ell)$ is not feasible. This naturally raises the question of identifying the simplest graph classes for which the problem remains computationally intractable. Dey et al. [8] proposed a polynomial-time algorithm for MCS on bi-colored paths, spiders, combs, etc. Also, Dey et al. [9] proposed a polynomial-time algorithm for bi-colored trees. Arimura et al. [10] introduced a XP algorithm parameterized by the number of colors $c$, with a running time of $\mathcal{O}(2^{4c}n^{2c+3})$. Recently, we also showed that MCS on trees is NP-complete in [11]. In the article [12], we demonstrate that fixed-parameter tractability has a faster running time for the same problem. Manna presented an $(4\alpha+2)$ approximation algorithm of MCS in interval graphs and APX-hard in circle graphs in the paper [13]. A recent paper [14] introduces the concept of the minimum consistent spanning subset, which is a variant of the consistent subset problem. The restriction is that every block (a block is a connected subgraph with the same color vertices) of the graph must contain at least one vertex in the consistent subset. In the paper [14], it is assumed that a consistent subset of a graph must contain at least one vertex from each block, despite the fact that in general, a consistent subset may not contain at least one vertex from each block. They have implemented algorithms in trees that require polynomial time due to these constraints.
@@ 2.2 New Results
We demonstrate that MSCS is NP-hard in general graphs in section 3. We show a 2-approximation in trees in the same section. Later, we show $\mathcal{O}(n^{4})$ polynomial-time algorithms in trees in the section 4 for finding MSCS. In section 5, we demonstrate faster polynomial-time algorithms in paths, spiders, and combs with running times $\mathcal{O}(n)$, $\mathcal{O}(n^{2})$, $\mathcal{O}(n^{3})$, respectively for finding MSCS.
@@ 3.1 Preliminaries
If all vertices of $T$ have the same color, i.e. $T$ is monochromatic, then every vertex of $T$ is a minimum strict consistent subset for $T$.
The definition of block as a maximal subset of connected vertices of the same color in $\mathrm{T}$. In view of Observation 3.0.1 we may assume that $\mathrm{T}$ has more than one block. Let $k\geq 2$ be the number of blocks in $\mathrm{T}$. We define the block tree of $\mathrm{T}$, denoted by $\mathcal{B}(\mathrm{T})$, as the tree with $k$ vertices, each corresponding to a block of $\mathrm{T}$, and there is an edge between any two vertices if their corresponding blocks are neighbors in $\mathrm{T}$. In other words, $\mathcal{B}(\mathrm{T})$ is obtained by contracting all blocks of $\mathrm{T}$. Notice that each vertex of $\mathcal{B}(\mathrm{T})$ corresponds to a block of $\mathrm{T}$, and vice versa. We refer to a block of $\mathrm{T}$ as a leaf block if its corresponding vertex in $\mathcal{B}(\mathrm{T})$ is a leaf. A block of $\mathrm{T}$ that is not a leaf block is called an internal block (see Figure 2(A) and 2(B)).
Figure 2: (A) Example of a tree $\mathrm{T}$ together with its blocks. (b) The block tree $\mathcal{B}(\mathrm{T})$ of $\mathrm{T}$.
@@
Every strict consistent subset has at least one vertex from each block of $T$.
@@
Any minimum strict consistent subset of a tree $T$ contains exactly one vertex from each leaf block of $T$.
@@
We use contradiction to prove this lemma. Consider a strict minimum consistent subset $S^{\prime}$ that contains more than one vertex from some leaf block, say $B$. Let $\mathrm{N}$ be the neighboring block of $B$. Let $x$ be a vertex of $S^{\prime}$ in $B$ that is closest to $\mathrm{N}$, as in Figure 3. Let $S^{\prime\prime}$ be the set obtained from $S^{\prime}$ by removing all the vertices of B except for $x$. We claim that $S^{\prime}$ is a strict consistent subset for $\mathrm{T}$; this would contradict $S^{\prime}$ being minimum.
To prove the above claim, it suffices to show that $x$ is a vertex of $S^{\prime}$ that is closest to every vertex of $B$. Let $t$ be a vertex of $S^{\prime}$ in $\mathrm{N}$ that is closest to $B$. Let $v$ be the last vertex of $B$ on the path from $x$ to $t$, as depicted in the figure to the right. As $S^{\prime}$ is a strict consistent subset, we have $\mathrm{d}(v,x)<\mathrm{d}(v,t)$. Now consider any vertex in $u\in B$. The path between $u$ and $t$ passes through $v$. Therefore $\mathrm{d}(u,x)<\mathrm{d}(u,t)$, which proves the claim. ∎
Figure 3: Illustration of Lemma 3.2.

### v3

@@ Abstract
Let $G$ be a simple connected graph with vertex set $V(G)$ and edge set $E(G$. Each vertex of $V(G)$ is colored by a color from the set of colors $\{c_{1},c_{2},\dots,c_{\alpha}\}$. We take a subset $S$ of $V(G)$, such that for every vertex $v$ in $V(G)\setminus S$, at least one vertex of the same color is present in its set of nearest neighbors in $S$. We refer to such an $S$ as a consistent subset (CS). The Minimum Consistent Subset (MCS) problem is the computation of a consistent subset of the minimum cardinality. It is established that MCS is NP-complete for general graphs, including planar graphs. The strict consistent subset is a variant of consistent subset problems. We take a subset $S^{\prime}$ of $V(G)$, such that for every vertex $v$ in $V(G)\setminus S^{\prime}$, all the vertices in its set of nearest neighbors in $S^{\prime}$ have the same color as that of $v$. We refer to such an $S^{\prime}$ as a strict consistent subset (SCS). The Minimum Strict Consistent Subset (MSCS) problem is the computation of a strict consistent subset of the minimum cardinality.
We demonstrate that MSCS is NP-hard for general graphs using a reduction from dominating set problems. We construct a 2-approximation algorithm and a polynomial-time algorithm in trees. Lastly, we conclude the faster polynomial-time algorithms in paths, spiders, and combs.
Keywords Minimum Strict Consistent Subset  $\cdot$ Paths  $\cdot$ Combs  $\cdot$ Spiders  $\cdot$ Trees  $\cdot$ Nearest-Neighbor  $\cdot$ Classification  $\cdot$ Approximation Algorithm  $\cdot$ NP-hardness.
@@ 1 Introduction
Many supervised learning approaches use a colored training dataset $T$ in a metric space $(X,d)$ as input, with each element $t\in T$ assigned a color $c_{i}$ from the set of colors $C=\{c_{1},c_{2},\dots,c_{\alpha}\}$. The purpose is to preprocess $T$ in order to meet specific optimality requirements by immediately assigning color to any uncolored element in $X$. The closest neighbor rule, which assigns a color to each uncolored element $x$ based on the colors of its $k$ closest neighbors in the training dataset $T$ (where $k$ is a fixed integer), is a popular optimality criterion. The efficiency of such an algorithm is determined by the size of the training dataset. As a result, it is necessary to reduce the cardinality of the training set while retaining the distance data. Hart [1] introduced this concept in $1968$ with the minimum consistent subset (in shortMCS) problem. Given a colored training dataset $T$, the goal is to find a subset $S\subseteq T$ of the minimum cardinality such that the color of every point $t\in T$ is the same as one of its nearest neighbors in $S$. Over $2800$ citations to [1] on Google Scholar demonstrate that the MCS problem has found several uses since its inception. [2] demonstrated that the MCS problem for points in $\Re^{2}$ under the Euclidean norm is NP-complete if at least three colors color the input points. Furthermore, it is NP-complete even with two colors [3]. Recently, it was established that the MCS problem is W[1]-hard when parameterized by the output size [4].[2] and [5] offer some other algorithms for the MCS problem in $\Re^{2}$.
1.1 Notations and Definitions
@@
This paper explores the minimum consistent subset problem when $(\mathcal{X},d)$ is a graph metric. Most graph theory symbols and notations are standard and come from [6]. For every graph $G$, we refer to the set of vertices as $V(G)$, and the set of edges as $E(G)$. Consider a graph $G$ with a vertex coloring function $C:V(G)\rightarrow\{c_{1},c_{2},\dots,c_{\alpha}\}$. For any subset $U$ of $V(G)$, let $C(U)$ represent the set of colors of the vertices in $U$. Formally, $C(U)=\{C(u):u\in U\}$. For any two vertices $u,v\in V(G)$, we use $\mathrm{d}(u,v)$ to signify the shortest path distance (number of edges) between $u$ and $v$ in $G$. $\mathrm{d}(v,U)=\min_{u\in U}\mathrm{d}(v,u)$ for a vertex $v\in V(G)$.
For any graph $G$ and any vertex $v\in V(G)$, let $\mathrm{N}(v)=\{u\mbox{ }|\mbox{ }u\in V(G),(u,v)\in E(G)\}$ indicate the set of neighbors of $v$ and $\mathrm{N}[v]=\{v\}\cup\mathrm{N}(v)$. For $U\subset V$, $\mathrm{N}[U]=\cup_{u\in U}\mathrm{N}[u]$. The distance between two subgraphs $G_{1}$ and $G_{2}$ in $G$ is represented as $\mathrm{d}(G_{1},G_{2})=\min\{\mathrm{d}(v_{1},v_{2})\mbox{ }|\mbox{ }v_{1}\in V(G_{1}),v_{2}\in V(G_{2})\}$. Therefore, $\mathrm{N}^{r}[v]=\{u\in V\mbox{ }|\mbox{ }\mathrm{d}(u,v)\leq r\}$. The nearest neighbor of $v$ in the set $U$ is indicated as $\mathrm{NN}(v,U)$, which is formally defined as $\{u\in U\mbox{ }|\mbox{ }\mathrm{d}(v,u)=\mathrm{d}(v,U)\}$. Also, for $X,U\subseteq V$, $\mathrm{NN}(X,U)=\cup_{v\in X}\{u\in U\mbox{ }|\mbox{ }\mathrm{d}(v,u)=\mathrm{d}(v,U)\}$. The shortest path in $G$ is $\mathrm{path}(v,u)$, which connects vertices $v$ and $u$. $\mathrm{path}(v,U)=\{\mathrm{path}(v,u)\mbox{ }|\mbox{ }u\in\mathrm{NN}(v,U)\}$. Without loss of generality, we shall use $[n]$ to denote the set of integers $\{1,\ldots,n\}$. A subset $X$ is said to be a consistent subset (in short CS) for $(G,C)$ if every $v\in V(G)$, $C(v)\in C(\mathrm{NN}(v,X))$. A subset $S$ is said to be a strict consistent subset (in short SCS) for $(G,C)$ if every $v\in V(G)$, $C(v)=C(\mathrm{NN}(v,S))$. MCS is a CS with minimum cardinality. Similarly, a minimum strict consistent subset (in short MSCS) refers to a strict consistent subset with minimum cardinality. The strict consistent subset problem in graphs is defined as follows:
@@
MSCS on Graphs Input: A graph $G$, a coloring function $C:V(G)\rightarrow\{c_{1},c_{2},\dots,c_{\alpha}\}$, and an integer $\ell$. Question: Does there exist a strict consistent subset of size $\leq\ell$ for $(G,C)$?
For a SCS $S$ and a vertex $v\in V(G)\setminus S$, $v$ is said to be covered by a vertex $u\in\mathrm{NN}(v,S)$ if $C(u)=C(v)=C(\mathrm{NN}(v,S))$. In other words, $u$ covers $v$. A block is a connected subgraph with the same color vertices. Suppose $k$ is the number of blocks in the tree $\mathrm{T}$. We define the block tree of $\mathrm{T}$, denoted by $\mathcal{B}(\mathrm{T})$, as the tree with $k$ vertices, each corresponding to a block of $\mathrm{T}$, and there is an edge between any two vertices if their corresponding blocks are neighbors in $\mathrm{T}$. In other words, $\mathcal{B}(\mathrm{T})$ is obtained by contracting all blocks of $\mathrm{T}$. Notice that each vertex of $\mathcal{B}(\mathrm{T})$ corresponds to a block of $\mathrm{T}$, and vice versa. We refer to a block of $\mathrm{T}$ as a leaf block if its corresponding vertex in $\mathcal{B}(\mathrm{T})$ is a leaf. A block of $\mathrm{T}$ that is not a leaf. The examples of blocks in a tree and the block tree are given in Figure 1(A) and 1(B), respectively. consistent spanning subset (in short CSS) is a variant of the consistent subset problem. A minimum consistent spanning subset (in short MCSS) is a minimum consistent subset with a restriction that every block of the graph must contain at least one vertex in the consistent subset. Note that not every block contains at least one vertex in the solution of MCS, but MCSS takes at least one vertex from each block in the solution. For example, we present MCS, MCSS and MSCS in Figure 2(A), 2(B) and 2(C), respectively (vertices inside the circle are in the solution). The examples above demonstrate that MCSS and MSCS are distinct variations of MCS. MSCS is an important version of MCS because all neighbor vertices of a vertex $v$ in the solution set must have the same color as $v$.
Figure 1: (A) Blocks are $B_{1},\dots B_{6}$. (B) The corresponding block tree.
@@
Figure 2: (A), (B) and (C) are the examples of MCS, MCSS and MSCS for a two-colored unweighted tree, respectively.
@@ 2 Previous Results and New Results
The previous results on graphs and our new results are given below.
@@ 2.1 Previous Results
Banerjee et al. [7] demonstrated that the MCS is  with when parameterized by the size of the minimum consistent set, even with only two colors. Under basic complexity-theoretic assumptions for generic networks, a FPT algorithm parameterized by $(c+\ell)$ is not feasible. This naturally raises the question of identifying the simplest graph classes for which the problem remains computationally intractable. Dey et al. [8] proposed a polynomial-time algorithm for MCS on bi-colored paths, spiders, combs, etc. Also, Dey et al. [9] proposed a polynomial-time algorithm for bi-colored trees. Arimura et al. [10] introduced a XP algorithm parameterized by the number of colors $c$, with a running time of $\mathcal{O}(2^{4c}n^{2c+3})$. Recently, Manna et al. also showed that MCS on trees is NP-complete in [11]. In the article [12], Manna et al. demonstrated that fixed-parameter tractability has a faster running time for the same problem. Also, Manna presented an $(4\alpha+2)$ approximation algorithm of MCS in interval graphs and APX-hard in circle graphs in the paper [13]. A recent paper [14] implemented a polynomial time algorithm for MCSS in trees. We took some basic ideas from the articles [14] and [8] to present the following new results.
@@ 2.2 New Results
We demonstrate that MSCS is NP-hard in general graphs in section 3. We show a 2-approximation in trees in the same section. However, the 2-approximation algorithm is not important because we show $\mathcal{O}(n^{4})$ polynomial-time algorithms for finding MSCS in trees in the section 4. In section 5, we demonstrate faster polynomial-time algorithms in paths, spiders, and combs with running times $\mathcal{O}(n)$, $\mathcal{O}(n^{2})$, $\mathcal{O}(n^{3})$, respectively for finding MSCS.
@@ 3 Preliminaries, NP-hardness in General Graphs and 2-Approximation Algorithm in Trees
Before discussing the algorithm for finding MSCS in trees, we are describing some basic properties as follows.
@@ 3.1 Preliminaries
If all vertices of $\mathrm{T}$ have the same color, i.e. $\mathrm{T}$ is monochromatic, then every vertex of $\mathrm{T}$ is a minimum strict consistent subset for $\mathrm{T}$.
@@
Every strict consistent subset has at least one vertex from each block of $\mathrm{T}$.
@@
As every block contains at least one vertex in the solution, so it is obvious that the vertices of $B$ must be covered, the vertices appear instead of taking $\alpha$ colors, we can put algorithms for only $2$ colors that must work for $\alpha$ colors.
@@
A minimum strict consistent subset of $\mathrm{T}$ includes exactly one vertex from each leaf block.
@@
We show this lemma by contradiction. Consider a minimum strict consistent subset $S^{\prime}$ with multiple vertices from leaf block $B$. Let $\mathrm{N}$ be the block next to $B$ (that is adjacent to $B$ in the block tree). Let $x$ be the nearest vertex of $S^{\prime}$ in $B$ to $\mathrm{N}$, as shown in Figure 3(A). The set $S^{\prime\prime}$ is obtained by eliminating all vertices of B except $x$ from $S^{\prime}$. We argue that $S^{\prime}$ is a strict consistent subset of $\mathrm{T}$, which contradicts that $S^{\prime}$ is minimum.
Now we show that $x$ is a vertex of $S^{\prime}$ that is closest to all vertex of $B$ to establish the above claim. Let $t$ be the nearest vertex of $S^{\prime}$ in $\mathrm{N}$ to $B$. The figure shows $v$ as the last vertex of $B$ on the path from $x$ to $t$. Since $S^{\prime}$ is a strict consistent subset, $\mathrm{d}(v,x)<\mathrm{d}(v,t)$. Now, examine any vertex in $u\in B$. The path from $u$ to $t$ travels through $v$. Thus, $\mathrm{d}(u,x)<\mathrm{d}(u,t)$, and our claim is proved. ∎
Figure 3: (A) Illustration of Lemma 3.2. (B) The reduced graph $G^{\prime}$.